OpenStudy (anonymous):

Solve 3c + 4 = -20. (1 point) -5.3 6 8 -8

OpenStudy (nurali):

3c + 4 = -20 3c=-24 c=-24/3 c=-8

OpenStudy (adilalvi):

-8

OpenStudy (anonymous):